OpenStudy (anonymous):

Express the statement as a formula. Use the given information to find the constant of proportionality, k. M is jointly proportional to a, b, and c, and inversely proportional to d. If a and d have the same value and if b and c are both 2, then M = 124.

OpenStudy (anonymous):

Alpha will be used as the proportionality symbol \[M \alpha \frac{ abc }{ d }\] This means that \[M = k \left( \frac{ abc }{ d } \right)\]You know the following as well \[a=d, b=2, c=2.\rightarrow M=124\]with the arrow signifying if the previous are true, then the following is true. Knowing this, solve for k.

OpenStudy (anonymous):

Let me know if you got it. If this helped you, feel free to give me a medal, and a testimonial if you feel i was of great assistance.

OpenStudy (anonymous):

solve for k please

OpenStudy (anonymous):

\[M = k \left( \frac{ abc }{ d }\right)\]plug in what you know \[124=k \left( \frac{ d ^{2} 2 }{ 2 } \right)\] the two's cancel out, leaving you with the following \[124=kd^2\]since d and a are interchangeable, you get either of the following answers \[k=\frac{ 124 }{ a^2 }\] or \[k=\frac{ 124 }{ d^2 }\] or \[k= \frac{ 124 }{ ad }\]

OpenStudy (anonymous):

none of those answers work it says that the answer cannot be understood

OpenStudy (anonymous):

it shouldnt be in a fraction form because it wont let me input fractions

OpenStudy (anonymous):

Well depends. Is there further information about A or B?

OpenStudy (anonymous):

A or D, i apologize.

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

Oh wow I do so humbly apologize! I read the initial equation incorrectly! Since a and d equal each other, they should cancel, and you are left with the following! \[M = k \left( bc \right)\] and since both b and c are equal to 2, you get the following \[M = k \left( 4 \right)\] and since M = 124, you get the following \[\frac{ 124 }{ 4 } =k\] \[k=31\]